### v3

@@ 7.3 Consideration of Sustainability
The imperative to reduce energy consumption in the context of AI and 6G integration is deeply rooted in the global consensus on carbon neutrality and peak carbon emissions [278]. Through agreements like the Paris Agreement, the international community has committed to significant environmental goals, including halving carbon emissions by 2030. The advanced capabilities of AI are crucial for managing the complex operations of 6G networks, which promise unprecedented speeds, low latency, and high reliability. However, the computational power required for AI comes at a high energy cost, which, if unchecked, could increase carbon emissions. This directly conflicts with the global movement towards reducing greenhouse gas emissions and achieving a sustainable, low-carbon future. Therefore, the concept of “Green” development has become a core tenet of 6G evolution.
Particularly, the energy consumption of AI is not only restricted to the communication network itself but also pertains to cloud computing centers and edge computing devices. For example, the computational requirements of LLM (such as GPT-4) and complex AI tasks (such as RL) have increased substantially [279]. In the training stage, it may consume several hundreds to thousands of MWh of energy, while the energy consumption in the inference stage is directly related to the invocation frequency, model scale, and real-time requirements. Meanwhile, the leasing costs of cloud computing resources are affected by computing power, usage duration, storage requirements, and market prices. Users typically bear these costs through pay-as-you-go, reserved instance, or subscription models and may indirectly pass them on to end consumers through subscription fees or service charges. Regarding the evaluation of energy efficiency, it can be analyzed from multiple aspects. The energy efficiency of the communication system itself is usually measured by indicators such as energy per bit or spectral efficiency. For the energy consumption assessment of cloud computing centers and the edge, a more comprehensive perspective is needed. For instance, the energy efficiency of a data center can be evaluated by the power usage effectiveness metric, which reflects the ratio of the total energy consumption of the data center to the energy consumption of IT equipment. The closer the power usage effectiveness is to 1, the higher the energy utilization efficiency.
To address these challenges, techniques such as pruning, quantization, and using neural architecture search can help create leaner models that maintain high accuracy while reducing computational load [280]. Developing energy-aware AI models that dynamically adjust their energy consumption based on the current network demands can contribute to a more sustainable operation [281]. Another aspect is optimizing data processing and transmission within 6G networks [282]. Leveraging AI for intelligent traffic routing and predictive maintenance can reduce unnecessary energy expenditure. Furthermore, deploying AI at the network’s edge can help localize data processing, reducing the energy costs associated with data transfer across long distances. Integrating renewable energy sources into 6G infrastructure is also critical to the green 6G vision. The telecommunications industry can significantly reduce its reliance on fossil fuels and its overall carbon footprint by powering data centers and network operations with solar, wind, and other renewable energies.
@@ 7.4 Requirement of Real-Time AI
The 6G aims to achieve uRLLC to support various real-time applications, including autonomous driving, telemedicine, and industrial automation. These scenarios impose unprecedented requirements on network real-time responsiveness and reliability, necessitating data transmission and processing to be completed within milliseconds or even microseconds. However, the current state-of-the-art AI models, characterized by large parameter spaces and intensive computation, typically employ distributed deployment, which inevitably introduces additional latency [283]. Furthermore, in data fusion scenarios that demand real-time data processing, traditional algorithms may encounter challenges in ensuring real-time performance. For example, in scenarios such as sensor data fusion in industrial automation and real-time traffic condition monitoring and analysis in intelligent transportation systems, algorithms are required to process data and make accurate fusion decisions within an extremely short time [284].
DL algorithms construct DNNs to automatically learn deep feature representations from vast and complex data, unlike traditional methods that rely on manually designed feature extraction rules. This significantly reduces errors and limitations caused by human factors, thereby enhancing the accuracy and real-time performance of data fusion. RL dynamically adjusts fusion weights and methods based on the real-time feedback information from different data sources to adapt to the constantly changing data environment and task requirements. Reducing the computational requirements of models without significantly compromising accuracy is possible through techniques such as model pruning, parameter quantization, and adopting more efficient network architectures, such as lightweight versions of CNN and RNN [285]. Additionally, adopting edge computing can reduce latency by shifting data processing tasks from the central cloud to the network edge [286]. Innovations in network architecture, such as network slicing technology, can also provide customized network resources for different services and applications, ensuring critical tasks receive the necessary bandwidth and latency guarantees [287]. Moreover, AI can manage network resources dynamically and accommodate real-time changes in network conditions. The development of hardware acceleration is also crucial for achieving real-time AI. Dedicated AI processors, such as GPU, Tensor Processing Units, and Field-Programmable Gate Arrays, can expedite the inference process of models, helping to meet the stringent real-time requirements of 6G networks.
@@ 7.5 Security and Privacy
The integration of AI and 6G introduces complex security and privacy challenges. AI algorithms are susceptible to adversarial attacks, where slight, often imperceptible, perturbations in the input data can lead to incorrect outputs, thereby compromising the system’s integrity. Moreover, the security and privacy challenges faced by AI algorithms in the process of data fusion are extremely formidable. Since the data being processed generally originates from numerous diverse data sources, which frequently encompass sensitive information such as personal identities, medical records, and financial transactions. This thereby gives rise to risks prevailing in all aspects of data fusion, ranging from the storage and transmission of data to its processing [288]. For instance, the server where the fused data is stored might be subject to hacker attacks, subsequently leading to data leakage. Moreover, during the data preprocessing stage of the AI models training, if appropriate encryption protection measures are lacking, the data is highly susceptible to being stolen.
Some existing work has inspired addressing these security and privacy issues. One aspect involves the development of robust AI models through techniques like adversarial training, which can enhance the resilience of AI systems against malicious to mitigate these inputs [289]. Another approach is the implementation of FL, where AI models are trained across multiple decentralized devices without exchanging the raw data, thereby preserving privacy [290]. Moreover, the authors in [291] proposed an adaptive feature-correlation region segmentation mechanism, which maintained a sound privacy protection for the DL model under a given moderate privacy budget. In addition to technological solutions, regulatory frameworks and policies must be established to define clear data protection and cybersecurity standards in the context of AI-6G integration. This includes setting guidelines for responsible AI usage, ensuring compliance with privacy regulations such as the general data protection regulation, and fostering international collaboration to combat cyber threats. Moreover, SMPC and homomorphic encryption can perform computations on encrypted data, ensuring data privacy even when processed by third parties [292].
@@ 7.6 Human-AI Interactive Collaboration
Integrating AI with 6G communication networks presents transformative opportunities for enhancing human-AI interactive collaboration. However, this convergence also introduces significant challenges that must be addressed to realize its full potential. One of the primary issues is the stringent real-time response requirements imposed by 6G’s ultra-low latency capabilities. conventional AI algorithms, which often require substantial computational resources and time, may struggle to develop this, so developing lightweight AI models optimized for edge computing is essential. These models must be capable of rapid processing and decision-making, reducing the dependency on centralized data centers and minimizing latency. Furthermore, 6G networks will facilitate the transmission of diverse data types, including video, audio, text, and sensor data, necessitating human-AI interactive collaboration systems to handle and interpret multimodal data adeptly. Advances in DL and multimodal learning technologies are crucial for effectively fusing and understanding these varied data streams, thereby enhancing the naturalness and intuitiveness of interactions.
In addition to technical advancements, the architectural design of AI deployment within 6G networks plays a pivotal role in optimizing human-AI interactive collaboration. The distributed AI architecture inherent to 6G offers new avenues for more flexible and efficient interaction experiences. By strategically deploying intelligent algorithms at the network edge and on terminal devices, achieving a more seamless and responsive human-AI interactive collaboration is possible. This distributed approach not only alleviates the computational burden on central servers but also enhances the adaptability and resilience of the system. Moreover, integrating AI and 6G necessitates robust security and privacy measures to protect sensitive user data, which is paramount in maintaining user trust and compliance with regulatory standards. Future research should focus on developing secure, privacy-preserving AI models and algorithms that operate efficiently within the 6G framework. By addressing these challenges, the integration of AI and 6G can significantly advance human-AI interactive collaboration, providing more intelligent, efficient, and user-friendly communication services that meet the evolving demands of future societies.
@@ 8 Future Works
As we look towards the future, integrating AI and 6G networks presents many research opportunities and challenges. The following areas highlight key directions for future work:
@@
The advent of 6G networks heralds a paradigm shift in managing network operations, with AI playing a central role. Adaptive learning mechanisms, leveraging AI, are set to become the cornerstone of smart network management. These mechanisms must handle the complexity and dynamism of 6G environments characterized by heterogeneous devices, variable traffic patterns, and the need for real-time processing. Research is expected to focus on developing ML models that can continuously learn and adapt without human intervention. These models should be scalable to manage the exponential increase in connected devices and robust to ensure consistent performance under varying network states. One promising approach is using DRL, which can dynamically optimize network decisions. DRL can adjust network parameters in real-time, ensuring optimal performance for changing network loads and user mobility patterns. Moreover, integrating transfer learning techniques can enable these AI models to apply knowledge gained in one domain to new but related problems. This is particularly relevant in 6G networks that will operate in diverse environments and must quickly adapt to new conditions. Transfer learning can reduce the need for extensive retraining of models, thereby saving computational resources and time. Another focus area can be the development of self-healing networks using AI. By incorporating predictive analytics, networks can anticipate failures and automatically reroute traffic or adjust parameters to mitigate the impact. This proactive approach to fault management will be critical for maintaining the high reliability and availability required by 6G applications.
Optimized Network Architectures
The underlying network must be optimized to cater to their requirements for AI applications to reach their full potential. 6G networks, with their promise of higher bandwidths and lower latencies, provide an ideal platform for AI applications but also pose significant challenges in terms of architectural design. Future research can delve into creating network architectures that are inherently flexible and can dynamically adapt to the needs of AI applications. Such architectures must support high data rates and the ability to process large volumes of data with minimal latency. AI can significantly enhance the capability of the network to manage these requirements by predicting traffic patterns and adjusting resources accordingly. A key component of such architectures is edge computing, which brings computational resources closer to the data sources. This is crucial for latency-sensitive AI applications that require near-instantaneous decision-making. Research should explore the development of edge computing solutions that can seamlessly integrate with central cloud resources, providing a balanced computational hierarchy. Furthermore, deploying AI at the network edge introduces new challenges in distributed learning and data privacy. The FL emerges as a viable solution, allowing AI models to be trained locally at the edge, with only model updates being shared to the central server. This reduces the bandwidth requirements and addresses privacy concerns by keeping sensitive data localized.
@@
The sustainability of 6G networks is a critical concern that must be addressed from an environmental and an economic perspective. AI has the potential to drive significant advancements in this area by optimizing network operations to reduce energy consumption. Future research should focus on developing AI models that can accurately predict the energy consumption of network components and optimize their operation to minimize waste. This includes the dynamic adjustment of network infrastructure, such as BSs and data centers, based on user demand and traffic patterns. AI can enable these components to enter low-power states when demand is low and quickly ramp up when the demand increases. In addition to optimizing existing infrastructure, research should explore the design of new, energy-efficient hardware that can support AI operations at the network edge. This includes the development of specialized processors that can run AI algorithms with high efficiency and low power consumption. Renewable energy sources in 6G networks can be another critical area of research. AI can manage integrating renewable energy sources, such as solar and wind, into the network’s power supply. This involves predicting energy generation patterns and adjusting network operations to match the availability of renewable energy. Furthermore, AI can contribute to developing new communication protocols designed with energy efficiency in mind. By optimizing how data is transmitted and received, these protocols can reduce the energy consumption of network devices and extend their operational lifetime.
@@
Ultra-low latency communications are a key feature of 6G networks, enabling a new generation of applications that require immediate response times. AI is poised to play a pivotal role in achieving this goal by providing real-time data processing and decision-making capabilities. Future research should focus on developing AI systems that can operate at the network edge, providing near-instantaneous response times for critical applications. This includes creating lightweight AI models that can be deployed on edge devices with limited computational resources. One of the challenges in this area is the need for AI algorithms that can make decisions based on incomplete or uncertain data. Research should explore the use of techniques such as probabilistic modeling and approximate computing to enable AI systems to operate effectively under these conditions. Another critical research area can be the development of AI-driven network orchestration tools that can manage the allocation of computational resources in real time. These tools must be able to predict the computational needs of different applications and allocate resources accordingly to ensure that latency requirements are met.
@@
The security of 6G networks is of utmost importance, as these networks will support a wide range of critical applications and services. AI has the potential to significantly enhance the cybersecurity of 6G networks by providing advanced threat detection and response capabilities. Future research can look into developing AI-based security frameworks that can identify and respond to cyber threats in real time. This involves using ML algorithms to analyze network traffic and detect anomalies that may indicate a security breach. One of the challenges in this area is the need for AI systems that can adapt to the constantly evolving landscape of cyber threats. Research should explore using adaptive learning algorithms that can update their models in response to new types of attacks. Another critical research area can be the development of AI-driven security protocols that can automatically patch vulnerabilities and respond to incidents without human intervention. This requires the creation of intelligent systems that can understand the context of security incidents and take appropriate actions to mitigate their impact. The use of AI for cybersecurity also raises significant privacy concerns. Research can address these concerns by developing privacy-preserving AI algorithms that protect user data while providing adequate security.
AI-Driven Innovations in 6G Communication Technologies
The AI with 6G communication networks promises significant advancements in key technologies such as intelligent reflecting surface (IRS) [293], integrated sensing and communication (ISAC), O-RAN, non-terrestrial networks (NTN) and near-field communication [294]. For example, IRS can dynamically manipulate wireless environments to optimize signal propagation. Future research should develop advanced AI algorithms for high dimensional and real-time IRS beamforming and channel estimation, leveraging ML to adapt to changing channel conditions in practical complex scenarios. The ISAC merges sensing and communication functions to improve spectrum utilization and provide environmental context for network optimization. AI-driven techniques for joint waveform design and resource allocation are crucial for achieving seamless ISAC integration [295]. Moreover, developing AI algorithms for real-time data fusion will be essential for extracting actionable insights from ISAC systems. ORAN’s software-defined architecture can benefit from AI-driven network management, enhancing adaptability and resilience. NTN, including satellite systems, can leverage AI for efficient resource management and fault detection, ensuring robust global connectivity. Future research should focus on AI models tailored to NTNs, uniquely addressing challenges like latency and dynamic network topology. By advancing AI applications in these key areas, researchers can unlock new levels of performance and efficiency in 6G networks, paving the way for a more connected and intelligent world.
@@
With the proliferation of intelligent devices and the advancement of IoT technologies, the demand for intelligent services from users and devices across wide areas continues to grow. Consequently, the integration of 6G and AI aims to provide seamless AI services to users and devices worldwide, achieving ubiquitous computing and meeting the diverse intelligent service needs of users and devices. By deploying edge servers on unmanned aerial vehicle (UAV) and low earth orbit (LEO) satellites to implement MEC, multi-layered and heterogeneous computing services can be delivered to users and devices, satisfying their differentiated QoS requirements. The primary challenge in this domain lies in designing rational, efficient, and dynamic resource allocation schemes for heterogeneous networks that integrate terrestrial BSs, UAVs, and LEO satellites. To address this, research should be based on the characteristics of integrated terrestrial, aerial, and satellite network architectures, leveraging theories such as convex optimization and Lyapunov optimization. Particular emphasis should be placed on advanced algorithms like DRL to explore computation offloading strategies and resource allocation scheme designs.

### v4

@@ 7.3 Consideration of Sustainability
While the advanced capabilities of AI are crucial for managing complex 6G operations to achieve unprecedented speeds, low latency, and high reliability, the computational power required comes at a high energy cost, which, if unchecked, could increase carbon emissions. The energy consumption of AI pertains to communication networks, cloud computing centers, and edge computing devices. For example, the computational requirements of LLM (e.g., GPT-4) and complex AI tasks ( e.g., RL/DRL) have increased substantially [314]. The training stage may consume several hundreds to thousands of MWh of energy, while the energy consumption in the inference stage is directly related to the invocation frequency, model scale, and real-time requirements [315]. For the energy consumption assessment of cloud computing centers and edge, a more comprehensive perspective is needed. For instance, the energy efficiency of a data center can be evaluated by the power usage effectiveness, which is the ratio of the total energy consumption of the data center to the energy consumption of IT equipment. The cost of leasing these computing resources is affected by the duration of resource usage, the required level of computing power, and market price fluctuations, and is ultimately passed on to end - users in the form of service pricing, subscription fees, etc.
Techniques such as pruning, quantization, and using neural architecture search can help create models that maintain high accuracy while reducing computational load [316]. Energy-aware AI models, which are capable of dynamically adjusting their energy consumption in response to network demands, can also contribute to sustainable operations [317]. Additionally, optimizing data processing and transmission within 6G networks represents another crucial aspect [318]. Integrating renewable energy sources into 6G infrastructure can also contribute to the green 6G vision.
@@ 7.4 Requirement of Real-Time AI
The 6G aims to achieve uRLLC in support of real-time applications, including autonomous driving, telemedicine, and industrial automation. These applications require data transmission and processing to be completed within milliseconds or even microseconds. However, the state-of-the-art AI models, characterized by large parameter spaces and intensive computation, typically employ distributed deployment and incur additional latency [319]. In scenarios such as sensor data fusion for industrial automation and traffic monitoring and analysis for intelligent transportation systems, algorithms are required to process data and make accurate fusion decisions within an extremely short time [320].
DL algorithms construct DNNs to extract deep feature representations from vast and complex data, unlike traditional methods that rely on manually designed feature extraction rules. This significantly reduces errors and limitations caused by human factors, thereby enhancing the accuracy and real-time performance of data fusion. RL/DRL dynamically adjusts fusion weights and methods based on real-time feedback information from different data sources, adapting to the constantly changing data environment and task requirements. Reducing the computational requirements of models without significantly compromising accuracy is possible through designing more efficient network architectures, e.g., lightweight versions of CNN and RNN [321]. Adopting edge computing can reduce latency by shifting data processing tasks from the central cloud to network edge [322]. Innovations in network architecture, e.g., network slicing [323], can also provide customized network resources for different services and applications, ensuring critical tasks receive the necessary bandwidth and latency guarantees. Additionally, dedicated AI processors, such as GPU, tensor processing units, and field-programmable gate arrays, can expedite model inference, helping meet the stringent real-time requirements of 6G networks.
@@ 7.5 Security and Privacy
AI models are susceptible to adversarial attacks, and imperceptible perturbations in the input data can lead to incorrect outputs [324]. Moreover, the security and privacy risks faced by AI algorithms in the process of data fusion are formidable. Specifically, the data generally originate from diverse data sources, which frequently encompass sensitive information, such as personal identities, medical records, and financial transactions. This gives rise to risks on all aspects of data fusion, ranging from storage and transmission to processing [325, 326]. For instance, the server storing the fused data might be subject to hacker attacks. During the data preprocessing of AI model training, data are susceptible to theft if appropriate encryption protection measures are not in place.
One defense approach is developing robust AI through adversarial training, which can enhance the resilience of AI systems [327]. Another approach is FL and/or MARL, where AI models are trained across multiple devices without exchanging the raw data, thereby preserving privacy [328, 329]. For example, the authors of [330] proposed an adaptive feature-correlation region segmentation mechanism to offer privacy protection for DL models under a given moderate privacy budget. SMPC and homomorphic encryption can also be performed to ensure data privacy even when some processing needs to be conducted at third parties [331]. Moreover, regulatory frameworks and policies must be established to define clear data protection and cybersecurity standards in the context of AI-6G integration. This includes setting guidelines for responsible AI usage, ensuring compliance with privacy regulations, e.g., general data protection regulation, and fostering international collaboration to combat cyber threats.
@@ 7.6 Human-AI Interactive Collaboration
The core challenges faced by human-AI interactive collaboration include trust, transparency, fairness, and ethics. The “black-box” nature of AI makes it difficult for humans to understand their decision-making logic, resulting in insufficient trust or overreliance. For instance, the complexity of DL models often makes it impossible for users to judge their reliability, thus affecting the effectiveness of collaboration [332]. Meanwhile, typically dynamic human-AI interactive collaboration requires high adaptability in task allocation and role switching. However, the existing collaboration frameworks are prone to unreasonable task allocation and delayed role switching [333]. Moreover, AI ethics and responsibility also need to be addressed urgently. For example, when an AI-driven decision is incorrect, it becomes crucial to establish a clear framework for assigning responsibility between humans and AI.
Through natural language explanations, visualization tools, and interactive interfaces, users can better understand the decision-making logic of AI and establish trust [334]. For example, the division of labor between humans and machines can be adjusted in real time according to task complexity and AI capabilities [335]. Coupled with an intuitive interface design, it can effectively reduce the difficulty for users to adapt to AI systems. For reducing bias and enhancing fairness, training with diverse and high-quality datasets and the application of fairness algorithms are crucial. Real-time bias-correction mechanisms can monitor and correct unfairness in AI-made decisions [336]. Furthermore, establishing a global ethical and legal framework will provide institutional guarantees for the development and application of AI. The boundaries of AI responsibility can be clarified through legislation and interdisciplinary ethical review mechanisms ensure the compliance of AI technology with social values and legal norms [337].
@@ 8 Future Works
This section delves into directions that revolve around 6G, with the aim of providing robust support for its innovations. In-depth research on these areas will help lay a foundation for the development of 6G.
@@
Adaptive learning mechanisms, leveraging AI, are set to become the cornerstone of smart network management. These mechanisms must handle the complexity and dynamism of 6G characterized by heterogeneous devices, variable traffic patterns, and the need for real-time processing. These models should be scalable to manage the exponential increase in connected devices and robust to ensure consistent performance under varying network states. One promising approach is DRL, which can dynamically optimize network decisions [268]. Integrating transfer learning techniques can enable AI models to apply knowledge gained in one domain to different yet relevant problems. Transfer learning can reduce the need for retraining models, saving computational resources and time. Another approach can be self-healing networks using AI. By incorporating predictive analytics, networks can anticipate failures and automatically reroute traffic or adjust parameters. This proactive approach to fault management will be critical for maintaining the high reliability and availability required by 6G applications.
Network Architecture Design
With their promise of higher bandwidths and lower latencies, 6G networks provide an ideal platform for AI applications but pose challenges in architectural design. Future research can delve into creating network architectures that are inherently flexible and can dynamically adapt to the needs of AI applications. On the other hand, AI can significantly enhance network capability to manage these requirements by predicting traffic patterns and adjusting resources. A key component of such architectures is edge computing, which brings computational resources closer to the data sources, which is crucial for latency-sensitive AI applications. Research is expected to develop edge computing solutions that can seamlessly integrate with central cloud resources. Furthermore, deploying AI at the network edge introduces new challenges in distributed learning and data privacy. FL emerges as a viable solution, allowing AI models to be trained locally at the edge, with only model updates being shared to the central server.
@@
The sustainability of 6G networks is critical environmentally and economically. AI has the potential to drive advancements in this area by optimizing network operations to reduce energy consumption. Future research can be devoted to AI models that can accurately predict the energy consumption of network components and optimize their operation to minimize energy use. This includes the dynamic adjustment of network infrastructure, such as BSs and data centers, based on user demand and traffic patterns. AI can enable these components to enter low-power states when demand is low and quickly ramp up when the demand increases. Future research can also explore the design of new energy-efficient hardware that can support AI operations at the network edge. This includes the development of specialized processors that can run AI algorithms with high efficiency and low power consumption. Moreover, AI can manage the integration of renewable energy sources, e.g., solar and wind, and wireless power transfer into the network’s power supply [338]. This involves predicting energy generation patterns and adjusting network operations to match the availability of renewable energy.
@@
Ultra-low latency communication is a key aspect of 6G networks [339]. AI is poised to play a pivotal role in achieving this goal by providing real-time data processing and decision-making capabilities. Future research is expected to develop AI systems that can operate at the network edge, providing near-instantaneous response time for critical applications. This includes creating lightweight AI models that can be deployed on edge devices with limited computational resources. One challenge is the need for AI algorithms that can make decisions based on incomplete or uncertain data. The use of techniques, such as probabilistic modeling and approximate computing, is a promising direction to enable AI systems to operate effectively under these conditions. Another challenge is the development of AI-driven network orchestration tools that can allocate computational resources in real-time. These tools are expected to predict the computational needs of applications and allocate resources accordingly to ensure latency.
@@
AI has the potential to enhance the cybersecurity of 6G networks by providing advanced threat detection and response capabilities [340]. Future research can look into developing AI-based security frameworks that can identify and respond to cyber threats in real-time. This involves using ML algorithms to analyze network traffic and detect anomalies that may indicate a security breach. One challenge is the need for AI systems that can adapt to the constantly evolving landscape of cyber threats. Efforts can be directed to explore adaptive learning algorithms that can update their models in response to new types of attacks. Another challenge is developing AI-driven security protocols that can automatically patch vulnerabilities and respond to incidents without human intervention [341]. This requires the creation of intelligent systems that can understand the context of security incidents and take appropriate actions.
AI-Driven 6G Innovations
The incorporation of AI within 6G networks can potentially drive technological innovations, including intelligent reflecting surface (IRS) [342], ISAC, O-RAN, non-terrestrial network (NTN), and near-field communication [343]. An IRS can dynamically augment wireless environments to optimize signal propagation, where AI can facilitate high-dimensional channel estimation and beamforming. ISAC merge sensing and communication functions to improve spectrum utilization and provide environmental context for network optimization, where AI can contribute to joint waveform design and resource allocation for seamless ISAC integration [344] and AI-based real-time data fusion can be essential to extract actionable insights from ISAC systems. ORAN can benefit from the AI-based network management of its software-defined architecture, enhancing adaptability and resilience. NTN, including satellite systems, can leverage AI for efficient resource management and fault detection. Interesting research directions can be AI models tailored to NTNs and addressing challenges like latency and dynamic network topology. By advancing AI applications in these novel areas, researchers can unlock new levels of performance and efficiency in 6G networks.
@@
With the proliferation of intelligent devices and the advancement of IoT technologies, the demand for intelligent services from users and devices continues to grow across wide areas. The integration of 6G and AI aims to provide AI services to users and devices worldwide, offering ubiquitous computing and meeting diverse intelligent service needs. By deploying edge servers on unmanned aerial vehicles (UAV) [345] or LEO satellites, multi-layered and heterogeneous computing services can be delivered, satisfying their differentiated QoS requirements. The primary challenge lies in designing rational, efficient, and dynamic resource allocation schemes for heterogeneous networks that integrate terrestrial BSs, UAVs, and LEO satellites. Compared to classical theories, such as convex optimization and Lyapunov optimization, particular emphasis needs to be placed on advanced algorithms like DRL to explore computation offloading strategies and resource allocation designs.
LLM-Driven Cognitive Networking for 6G
The rise of low-cost open-source LLMs, such as DeepSeek [346], is set to drastically lower barriers to AI-driven innovation. These accessible frameworks enable global researchers and organizations to rapidly prototype and deploy lightweight LLMs for 6G’s mission-critical applications. The integration of LLMs with 6G networks is poised to revolutionize cognitive networking by enabling adaptive resource allocation [347], intent-based automation [348], embodied AI [349], and semantic communication [350]. Future research aims to embed LLMs into edge devices through lightweight architectures for real-time network optimization, while addressing challenges such as computational overhead and privacy risks. Key directions include federated learning for distributed LLM training, cross-modal semantic encoding to reduce bandwidth consumption, and DT-aided network simulation. Innovations in sparsity-aware inference and hybrid quantum-classical frameworks may unlock scalable deployment across 6G infrastructures. However, balancing model compression with performance retention and ensuring robust decision-making under dynamic conditions remain critical hurdles.